1. A 1.00 liter solution contains 0.46 M nitrous acid and 0.35 M sodium nitrite. If...

1.

A 1.00 liter solution contains 0.46 M nitrous acid and 0.35 M sodium nitrite.

If 30.0 mL of water are added to this system, indicate whether the following statements are true or false.

(Note the the volume MUST CHANGE upon the addition of water.)

_______TrueFalseA. The concentration of HNO2 will remain the same.

_______TrueFalseB. The concentration of NO2- will decrease.

_______TrueFalseC. The equilibrium concentration of H3O+ will remain the same.

_______TrueFalseD. The pH will increase.

_______TrueFalseE. The ratio of [HNO2] / [NO2-] will decrease.
Answer #1

Statement A: False

Explanation: Concentration of HNO2 = 0.46 mol/1.03 L = 0.4466 M (decreased).

Statement B: True

Explanation: Concentration of NO2- = 0.35 mol/1.03 L = 0.3398 M (decreased).​

Statement C: True

Statement D: False

Explanation: pH = -Log[H+], will be the same since [H+] is the same.

Statement E: False

Explanation: Initial ratio = 0.46/0.35 = 1.3143

Ratio after adding water = 0.4466/0.3398 = 1.3143

Both are the same.